OpenStudy (anonymous):

Slope of -15/5 passing through the points (-2.5, 4.5) what is the y intercept?

OpenStudy (anonymous):

use the point slope form y-y1=m(x-x1) y-4.5=-15/5(x-(-2.5)) y-4.5=(-15/5)x-7.5 y=(-15/5)x-3 so y intercept would be -3

OpenStudy (anonymous):

Thank ya for the help!

OpenStudy (anonymous):

no prob :)